OpenStudy (anonymous):

3.Which of the following points is a solution to y ≤ −2x + 3? A. (0, 4) B. (1, 0) C. (3, −2) D. (5, −2)

OpenStudy (anonymous):

@nephycx

OpenStudy (anonymous):

@123AB456C

OpenStudy (123ab456c):

hi!!:):)

OpenStudy (anonymous):

helpme please @123AB456C

OpenStudy (anonymous):

@jazzygirl127

563blackghost (563blackghost):

You can substitute.... So we can input the points for answer A....\[4\le-2(0)+3\]We need to make sure that our total on the right side is great than or equal to 4...

OpenStudy (anonymous):

how does this work

563blackghost (563blackghost):

B. \[0 \le -2(1)+3\] C. \[-2\le-2(3)+3\] D.\[-2 \le -2(5)+3\]

OpenStudy (cowgirl1):

the answer is Bjust plug in the x nad y values an see if they make the inequality true :-)

OpenStudy (jdoe0001):

well.. have you tested the given choices yet?

OpenStudy (anonymous):

I think it b or a

563blackghost (563blackghost):

@cowgirl1 you can not give away the answer to the openstudier it is against the code of conduct....

OpenStudy (jdoe0001):

b or a? how so?

OpenStudy (anonymous):

wait its not b or a I did it on my paper I think its c

OpenStudy (cowgirl1):

sorry I ddint know..i saw replies that way and thougt it was ok

OpenStudy (jdoe0001):

c? how did you get c?

563blackghost (563blackghost):

@cowgirl1 if you see replies where the answer is given please tell them not to otherwise they can be reported....

OpenStudy (jdoe0001):

@cowgirl1 Give a man a fish and you feed him for a day. Teach him how to fish and you feed him for a lifetime. ~~ Lao Tzu ~~

563blackghost (563blackghost):

@jdoe0001 wow I can see how that is related :)....if heard of that quote before ^^

563blackghost (563blackghost):

i've *

OpenStudy (cowgirl1):

ive seen it too...and again im sorry

OpenStudy (anonymous):

can some please explain me how to do this im cofused

OpenStudy (jdoe0001):

gimme a sec... lemme show you using A) choice

OpenStudy (cowgirl1):

ok on one side you have y and on the other you have x. the coordinates given are (x,y) if you fill in the values of x and y from the coordinates given into the inequality, you can check to see if the inequality is true

563blackghost (563blackghost):

You would use the points and substitute where correspond with.... the points are set like this (x,y) so we would input the numbers where each go.... ex. You are given (0,3) You have the eqaution y<2x+4 You would input 3 where y is and 0 where x is... 3<2(0)+4

563blackghost (563blackghost):

And its alright @cowgirl1 ^^

OpenStudy (jdoe0001):

hmmm hold the mayo a sec....lemme fix that.. I got the letters swapped shoot

OpenStudy (cowgirl1):

and thank you @563blackghost ...i wont do that again

OpenStudy (jdoe0001):

\({\color{brown}{ y}}\le-2{\color{blue}{ x}}+3\qquad \qquad \begin{array}{llll} A)&0&,&4\\ &x&&y \end{array} \\ \quad \\ \quad \\ {\color{brown}{ 4 }}\le-2{\color{blue}{ (0)}}+3\implies 0\le 0+3\implies 4\le3\) so.. is that really true? is \(4\le 3?\)

OpenStudy (jdoe0001):

so.. what do you think @gigiheart63 ... is that true? is really \(\large 4\le 3?\)

OpenStudy (jdoe0001):

shoot....I have a typo... but the result is correct... anyhow \({\color{brown}{ y}}\le-2{\color{blue}{ x}}+3\qquad \qquad \begin{array}{llll} A)&0&,&4\\ &x&&y \end{array} \\ \quad \\ \quad \\ {\color{brown}{ 4 }}\le-2{\color{blue}{ (0)}}+3\implies 4\le 0+3\implies 4\le3\)

OpenStudy (anonymous):

yes idk

OpenStudy (jdoe0001):

yes? idk? which is it? lemme put it this way is 4 Greater, or bigger than 3? if you have 4 oranges, as opposed to 3 oranges is 4 more than 3?

OpenStudy (jdoe0001):

or is 3 more than 4 or 4 is less than 3? would you rather have 3 oranges instead?

OpenStudy (jdoe0001):

or would you rather have avocados and chips =)

OpenStudy (anonymous):

4 is greter its its false

OpenStudy (jdoe0001):

ahah.... see... now we know A) choice is false, because, yes, you're correct 4 is "more" or greater than 3 thus \(\large 4\bcancel{\le} 3\) so... now.. .try the others... say B) \({\color{brown}{ y}}\le-2{\color{blue}{ x}}+3\qquad \qquad \begin{array}{llll} B)&1&,&0\\ &x&&y \end{array} \\ \quad \\ \quad \\ {\color{brown}{ 0 }}\le-2{\color{blue}{ (1)}}+3\implies ?\)

OpenStudy (jdoe0001):

anyway... check the given choices... see which one gives you true :)

563blackghost (563blackghost):

please close your question once you have gotten your answer ^^

alones (alones):

Wait so you got your answer answered???

OpenStudy (mathmale):

By far the simplest way to identify a solution to \[y \le-2x+3\]

OpenStudy (mathmale):

is to substitute the coordinates of each given point and determine whether the resulting inequality is true or false. A(0,4) has x=0 and y=4. Is this true or false? \[4\le-2(0)+3\] If true, then A is a solution; if false, then A is not a solution. Do the same thing in the case of the other 3 possible solutions.